A book containing 150 pages has 100 misprints. Find the probability that a particular

page contains (a) no misprints, (b) 5 misprints, (c) at least 2 misprints, (d) more than 1

misprint.


1
Expert's answer
2021-09-14T06:05:14-0400

Let "X=" the number of misprints per page: "X\sim Po(\lambda)."

"\lambda=\dfrac{100}{150}=\dfrac{2}{3}"


(a)


"P(X=0)=\dfrac{e^{-2/3}(\dfrac{2}{3})^0}{0!}=e^{-2/3}\approx0.513417"



(b)


"P(X=5)=\dfrac{e^{-2/3}(\dfrac{2}{3})^5}{5!}\approx0.000563"

(c)


"P(X\geq2)=1-P(X=0)-P(X=1)"

"=1-\dfrac{e^{-2/3}(\dfrac{2}{3})^0}{0!}-\dfrac{e^{-2/3}(\dfrac{2}{3})^1}{1!}\approx0.144305"


(d)


"P(X>1)=P(X\geq2)=1-P(X=0)-P(X=1)"

"=1-\dfrac{e^{-2/3}(\dfrac{2}{3})^0}{0!}-\dfrac{e^{-2/3}(\dfrac{2}{3})^1}{1!}\approx0.144305"